OpenStudy (anonymous):

Jake scored about twice as many points as wille jake scored 22 points write an equation and use it to estimate the number n of points willie scored.

OpenStudy (anonymous):

22 x 2 = n

OpenStudy (anonymous):

44?

OpenStudy (anonymous):

oh!!! the sentence would actually be 22 / 2 = n

OpenStudy (anonymous):

A n + 2 = 22 about 20 points B n - 2 = 22 about 24 points C 2n = 22 about 11 points D n/2 = 22 about 44 points

OpenStudy (anonymous):

c

OpenStudy (anonymous):

are u sure